1.679 grams of potassium hydrogen phthalate (molar mass 204.22 g/mol) is dissolved in 35.64 mL of distilled water. The potassium hydrogen phthalate solution is titrated and completely neutralized with 16.83 mL NaOH. What is the molarity of the NaOH solution?
1
Expert's answer
2020-09-29T06:21:52-0400

the number of moles:

mass of KHP = 1.679g

Molar mass of KHP = 204.22 g/mol

"n=\frac{m}{M}=\frac{1.679}{204.22}=0.008"

1 mole of KHP reacts with 1 mole of NaOH.

KHP + NaOH ==> NaKP + H2O

So, 0.008 moles of KHP will react with =  of KOH.

To calculate the molarity of KOH, we use the equation:

Moles of KOH = 0.008 moles

Volume of solution = 16.83 mL = 0.01683 L   1L = 1000 mL

according to formula^

"C=\frac{n}{V}=\frac{0.008}{0.01683}=0.489"

molarity of NaOH solution is 0.489 M.